# L4P: Low-Level 4D Vision Perception Unified

### v2

@@ Abstract
The spatio-temporal relationship between the pixels of a video carries critical information for low-level 4D perception tasks. A single model that reasons about it should be able to solve several such tasks well. Yet, most state-of-the-art methods rely on architectures specialized for the task at hand. We present L4P, a feedforward, general-purpose architecture that solves low-level 4D perception tasks in a unified framework. L4P leverages a pre-trained ViT-based video encoder and combines it with per-task heads that are lightweight and therefore do not require extensive training. Despite its general and feedforward formulation, our method matches or surpasses the performance of existing specialized methods on both dense tasks, such as depth or optical flow estimation, and sparse tasks, such as 2D/3D tracking. Moreover, it solves all tasks at once in a time comparable to that of single-task methods.
@@ 1 Introduction
Large collections of videos are our most complete and compact source of priors about the world. Much like text did for large-language models, the corpus of videos we amassed over the years allowed video-language models (VLMs) to produce remarkable zero-shot results on high-level vision tasks such as video captioning, video question answering, and others. However, zero-shot, low-level 3D and 4D vision perception tasks, such as depth from video, tracking, optical flow, and others remain a challenge. Pre-trained diffusion models fine-tuned on target-domain data showed potential on dense vision perception tasks (e.g., depth [28, 21], flow [44], etc.), but the fine-tuning makes them task-specific, and therefore limits their ability to leverage priors across multiple tasks at once. Sparse vision perception tasks, such as tracking, are even more challenging to tackle with a general foundation model, because their representation does not fit naturally into data structures like dense pixel planes. As a result, they have been historically addressed with carefully designed approaches, which are often optimization-based [63], or with specialized architectures [27, 8].
@@
This goal presents multiple challenges. To learn and share priors across tasks—including those beyond 4D perception—we need a strong backbone shared across tasks. The architecture should also be general to allow for pre-training on auxiliary tasks. Not least, dense and sparse tasks require fundamentally different representations, e.g., dense 2D planes vs. 3D tracks.
@@
Figure 1: We propose L4P, a general-purpose architecture that solves several low-level 4D perception tasks. Building on a pre-trained video encoder and lightweight per-task heads, our unified model matches or surpasses the performance of SOTA methods specialized to solve individual tasks. L4P can easily be extended to additional tasks, (e) and (f), without compromising performance.
@@
We tackle these challenges by combining a pre-trained video masked auto-encoder (VideoMAE) [59, 62] with per-task, lightweight heads (Figure 2). VideoMAEs have been successfully employed for a variety of mid- and high-level vision tasks [62], but their ability to capture the spatio-temporal relationship between pixels is underexplored in the context of low-level 4D perception. VideoMAEs offer a feedforward, online mechanism to tokenize videos within a small computational envelope. For dense tasks, we couple VideoMAE with heads based on the dense prediction transformer (DPT), which has been shown to perform well on depth estimation, image segmentation, and others [42]. For sparse tasks we focus on tracking, and specifically, on the track-any-point (TAP) formulation [7]. We posit that tracking is important for perception because understanding fine-grained, complex motions and the physical interaction between objects is critical to downstream applications, including 3D reconstruction [33, 61] and robotics manipulation [76, 72, 73, 1]. However, the point tracking tasks have historically been addressed using specialized architectures [27, 8, 74]. To unify them with other dense tasks, we formulate the problem of estimating tracks as that of predicting, for queried pixels, 2D heatmaps with associated depth and visibility tokens. This is the mechanism that allows us to tackle sparse and dense tasks within a unified model. To implement it, we augment a general-purpose head with a novel memory mechanism to track points for arbitrarily long videos. Thanks to the training curriculum we propose, and to VideoMAE’s priors, our unified model finetuned for multiple tasks at once on a small collection of synthetic datasets shows strong generalization to real data.
@@
Our formulation presents several desirable properties and advantages. First, the pre-trained VideoMAE model allows us to tap into priors learned from large datasets that are potentially different and more varied than those typically used for low-level 4D perception. It also affords us efficient computation: our system solves all tasks in roughly 300ms for a 16-frame video chunk ($\sim$19ms/frame), which is comparable to, or faster than, methods specialized for each task (see Table 1 in the Supplementary). Moreover, combining it with per-task heads allows us to train a relatively small number of parameters for new tasks, which we show by freezing the system and adding a head for motion-based segmentation and camera pose estimation (marked in green in Figures 1 and 2). Lastly, but perhaps most importantly, breaking the architecture into a general VideoMAE and per-task heads offers a mechanism to solve both dense and sparse tasks with a unified model (Figure 1). Despite being general, our architecture performs on par with, or better than, state-of-the-art methods. This is remarkable because the baselines we compare against are task-specific, carefully designed for, and specialized to excel at their respective tasks. Finally, VideoMAEs are already used as encoders for VLMs [70, 34], and we speculate that training them to reason about low-level 4D perception may impart those capabilities to the downstream VLMs they may be used with, though validating this statement is outside the scope of our paper.
@@ 2.1 Foundation models for vision perception
Self-supervised pre-training of large models on huge unlabeled data has shown great success. Among them, Vision Transformers (ViT) [30] pre-trained with masked autoencoding (MAE) [20] have become a common choice to fine-tune for many vision tasks, Segment Anything (SAM) [29] being a notable example. VideoMAE [59] and MAE-ST [12] adopt MAE on videos for spatio-temporal representation learning. VideoMAEv2 [62] introduces a dual masking strategy that allows them to efficiently scale up the model to a billion parameters. Their representation is effective for action recognition, and more recently is combined with language models [70, 34]. However, their uses remain unexplored for the low-level 4D perception tasks, which we address in this work.
@@ 2.2 Dense Prediction Tasks
Depth Estimation. Stereo approaches [14, 15, 47, 80, 22, 45] estimate depth by estimating correspondences between posed images, which limit their uses to static scenes. Depth estimation from a single image [11, 41, 42, 78, 79] has made significant progress, but suffers from 3D and temporal inconsistency in a video setting. We focus on estimating depth from videos of dynamic scenes, for which several solutions exist [35, 31, 54, 69]. To tackle the data scarcity problem, the state-of-the-art approaches leverage priors from foundation models: DepthAnything [78, 79] uses DINOv2 [38], DUSt3R [65] uses cross-view completion pre-training [71], and Marigold [28] and DepthCrafter [21] use image and video diffusion priors respectively. We benefit from large-scale pre-training by using a video encoder pre-trained with MAE.
@@
Camera pose estimation. Classical structure-from-motion solutions [46, 51] work robustly and accurately in constrained settings, while learning-based solutions [67, 56, 65, 60, 81, 64] address challenging scenarios like limited image observations, dynamic and texture-less regions, etc. We cast camera pose estimation as a dense task by representing cameras as dense bundles of rays [17, 82] using 6-D Plücker coordinates [40] and estimating them using our dense head.
@@ 2.3 Sparse Prediction Tasks
Tracking Any Point (TAP) in a video has many applications [73, 1, 33, 61, 6]. Particle Video [19], PIPs [3] and TAP-Net [7] lay the initial foundation by adopting ideas from optical flow approaches. On the other hand, OmniMotion [63] optimizes a volumetric representation for each video to solve this task and is time-consuming. TAPIR [8] introduces the idea of coarse-to-fine track estimation, BootsTAPIR [9] further improves it by adopting a self-supervised learning, and CoTracker [27] proposes to jointly track multiple points to leverage spatial correlations. SpaTracker [74] is one of the first feedforward 3D point tracking approaches. It uses estimated depth maps to uplift pixels to 3D and leverages CoTracker’s tracking formulation to track points in 3D.
Unlike previous approaches with specialized architectures, our solution solves point tracking in a unified way with other tasks. We share some motivations with Tracking at Any Granularity [18], a concurrent work. While they focus on training a large model from scratch on a large collection of datasets for tracking various types of queries, we choose to leverage a large-scale pre-trained model that lends itself to a wide range of tasks.
@@ 3.2 Dense Prediction Heads
For videos longer than $T$ frames, we run inference with stride $T/2$ and enforce consistency in the overlapping frames between temporal windows. For depth, we align the predictions with an affine transformation. This strategy has no effect on the individual windows for relative depth, but can greatly improve long-term temporal consistency. For optical flow and motion-based segmentation, we simply overwrite the predictions in the overlap between windows. For pose estimation, we align the camera trajectories in the overlapping windows with a similarity transformation.
Figure 4: We show results for perception tasks for scenes with both camera and object motion.
@@ 3.3.2 A Memory Mechanism for Long Videos
To track beyond a single window of length $T$ frames, we propose an online strategy. A naïve approach would be to chain tracks across windows. Given two consecutive and overlapping windows, and a 2D track estimated in the first one, we can use a point on the track in the temporal overlap between the two windows as the query for tracking in the second one. To pick a good point to chain the tracks, we can select the one with highest visibility score. However, this solution is brittle for two reasons. First, a tracked point may not be visible in the overlap between the windows. To tackle this problem, inspired by Karaev et al. [27], we introduce a track-feature token $\mathcal{F}$ that is passed to subsequent windows as an additional prompt (see Figure 3). However, unlike Karaev et al., we do not initialize it explicitly with the local appearance around the query point, so the two-way attention head is free to capture the most useful information to track through occlusions. Second, the naïve solution described above does not allow the system to reason across temporal windows, which makes it prone to drifting or to losing tracks. The track-feature tokens help, but to provide even more cross-window information, we pass the video tokens decoded by the two-way attention stage of the current window to the next, as shown in Figure 3. We achieve this by projecting the decoded video tokens in the overlapping region via a linear layer, and by adding them to the corresponding incoming video tokens to the two-way attention stage in the next window. Our memory strategy based on these two mechanisms is critical to allow proper reasoning across temporal windows as shown by the comparison in ablation study in Section 5.4.
@@ 4 Implementation
Training datasets. We use the video encoder from Wang et al. [62], which is pre-trained on 1.35M video clips for masked auto-encoding. To fine-tune our model, we use four synthetic datasets covering various types of annotations, and rely on the priors from the pre-trained video encoder for generalization. We use Kubric [16], a synthetic dataset in which multiple objects interact, annotated with ground truth (GT) depth, flow, and 2D/3D point tracking. To include videos with long 3D trajectories, we add PointOdyssey [83] and DynamicReplica [26]. Both are synthetic datasets with animated characters in mostly indoor scenes. Both have depth GT and DynamicReplica also offers optical flow GT. To further increase scene diversity, we also include TartanAir [68], which provides GT for flow and depth.
@@ 5 Experiments
Here we discuss our results and ablate the contribution of different components of our approach. We also show that L4P can be extended to new tasks with no performance degradation for the existing ones.
@@
AbsRel $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ $\delta_{1}$uparrow
@@
Table 1: Depth estimation results. We compare our full model with single-image (Row 1-3) and video depth estimation baselines (Row 4-6). On video datasets (all except NYUv2), our model consistently performs better than DepthCrafter, the closest competition, and by a large margin on ScanNet and KITTI.
Figure 5: Comparisons with SOTA methods for different tasks. We run inference on video clips and show only 1 frame for the top three tasks. Purple overlays and dotted trajectories represent GT for motion-based segmentation and camera pose estimation.
@@ 5.1 Video Depth Estimation
We follow DepthCrafter [21] and evaluate video depth estimation on a collection of five datasets. We do not use any of the datasets for training our models or the baselines to better understand their generalization abilities. When estimating depth for videos longer than $T=16$ frames, we follow the online inference and alignment strategies explained in Section 3.2. There is an inherent scale-ambiguity in the estimated depthmaps. We follow the common practice of aligning linearly the estimation with the GT before calculating evaluation metrics. The alignment is done for all the frames at once, and is carried out in disparity space via least-square fitting. For comparison on single image datasets, we repeat the single frame 16 times to compute our estimations. We report two metrics: AbsRel ($\text{mean}(|\hat{\mathbf{d}}-\mathbf{d}|/\mathbf{d}))$ and $\delta_{1}$ (ratio of pixels satisfying $\max(\mathbf{d}/\hat{\mathbf{d}},\hat{\mathbf{d}}/\mathbf{d})<1.25$), where $\mathbf{d}$ represents GT, and $\hat{\mathbf{d}}$ is depth estimation after alignment. We upsample our estimations from $224\times 224$ to each dataset’s original resolution for evaluation.
We consider video approaches including NVDS [69], ChronoDepth [49], DepthCrafter [21], as well as single-image ones, including Marigold [28] and DepthAnything [78, 79]. Among them, DepthCrafter [21] and DepthAnything [78, 79] each represent the SOTA respectively. Marigold and DepthCrafter afford impressive levels of details, but their denoising process is expensive.
Our results show consistent advantages over both SOTA single-image and video depth approaches on the four video datasets (Table 1). Since L4P is a video approach, applying it on single images from NYUv2 does not provide the necessary temporal context for it to perform well. DepthCrafter also similarly suffers on NYUv2. Figure 5 (first row) shows qualitative samples and comparison with select SOTA approaches. L4P produces a level of details comparable to that of diffusion models such as DepthCrafter, while generally capturing more accurate relative scales. It is also worth noting that our approach performs competitively on KITTI, despite the fine-tuning not including driving scenarios.
Kubric Dynamic Replica Spring
$EPE\downarrow$ $EPE<1\uparrow$ $EPE\downarrow$ $EPE<1\uparrow$ $EPE\downarrow$ $EPE<1\uparrow$
RAFT* [55] $0.31$ $94.6$ $0.14$ $98.7$ $0.13$ $98.4$
MemFlow [10] $0.27$ $95.6$ $0.11$ $99.2$ $0.13$ $98.4$
L4P $(0.13)$ $(97.5)$ $(0.03)$ $(99.9)$ $\mathbf{0.09}$ $\mathbf{98.7}$
Table 2: Optical flow estimation results. ($\cdot$) indicates that our method was trained on the training set of that dataset. On Spring, we surpass both two-frame (marked with *) and multi-frame SOTA baselines, despite the latter having a specialized architecture and an advanced memory mechanism.

### v3

# L4P: Towards Unified Low-Level 4D Vision Perception

@@ Abstract
The spatio-temporal relationship between the pixels of a video carries critical information for low-level 4D perception tasks. A single model that reasons about it should be able to solve several such tasks well. Yet, most state-of-the-art methods rely on architectures specialized for the task at hand. We present L4P, a feedforward, general-purpose architecture that solves low-level 4D perception tasks in a unified framework. L4P leverages a pre-trained ViT-based video encoder and combines it with per-task heads that are lightweight and therefore do not require extensive training. Despite its general and feedforward formulation, our method is competitive with existing specialized methods on both dense tasks, such as depth or optical flow estimation, and sparse tasks, such as 2D/3D tracking. Moreover, it solves all tasks at once in a time comparable to that of single-task methods.
@@ 1 Introduction
Large collections of videos are our most complete and compact source of priors about the world. Much like text did for large-language models, the corpus of videos we amassed over the years allowed video-language models (VLMs) to produce remarkable zero-shot results on high-level vision tasks such as video captioning, video question answering, and others. However, zero-shot, low-level 3D and 4D vision perception tasks, such as depth from video, tracking, optical flow, and others remain a challenge. Pre-trained diffusion models fine-tuned on target-domain data showed potential on dense vision perception tasks (e.g., depth [31, 24], flow [49], etc.), but the fine-tuning makes them task-specific, and therefore limits their ability to leverage priors across multiple tasks at once. Sparse vision perception tasks, such as tracking, are even more challenging to tackle with a general foundation model, because their representation does not fit naturally into pixel-aligned data structures, and are typically addressed with optimization-based approaches [69] or specialized architectures [30, 12].
@@
Figure 1: We propose L4P, a general-purpose architecture that solves several low-level 4D perception tasks. Building on a pre-trained video encoder and lightweight per-task heads, our unified model is competitive with existing methods specialized to solve individual tasks. L4P can easily be extended to additional tasks, (e) and (f), without compromising performance.
@@
This is challenging: we need a shared backbone that is strong and versatile. On one hand, it must be general enough to allow for pre-training on auxiliary tasks. On the other hand, it must be able to support diverse tasks that require fundamentally different representations, such as dense 2D planes and 3D tracks. We tackle this challenge by combining a pre-trained video masked auto-encoder (VideoMAE) [64, 68] with per-task, lightweight heads (Figure 2). VideoMAEs offer a feedforward, online mechanism to tokenize videos within a small computational envelope. Indeed they have been successfully employed for a variety of mid- and high-level vision tasks [68], but their ability to capture the spatio-temporal relationship between pixels is underexplored in the context of low-level 4D perception. For dense tasks, we couple VideoMAE with heads based on the dense prediction transformer (DPT), which has been shown to perform well on depth estimation, image segmentation, and others [47]. For sparse tasks, we focus on tracking. We posit that tracking is important for perception because understanding fine-grained, complex motions and the physical interaction between objects is critical to downstream applications, including 3D reconstruction [36, 66] and robotics manipulation [83, 78, 79, 2]. We formulate the problem of tracking as that of predicting 2D heatmaps for queried pixels with associated depth and visibility tokens. This formulation allows us to tackle sparse and dense tasks within a unified model. To implement it, we augment a general-purpose head with a novel memory mechanism to track points for arbitrarily long videos. We finetune our model for multiple tasks at once on a small collection of synthetic datasets. Yet, thanks to the training curriculum we propose, and to VideoMAE’s priors, it shows strong generalization to real data.
@@
Our formulation presents several desirable properties and advantages. First, the pre-trained VideoMAE model allows us to tap into priors learned from large datasets that are potentially different and more varied than those typically used for low-level 4D perception. It also affords us efficient computation: our system solves all tasks in roughly 300ms for a 16-frame video chunk ($\sim$19ms/frame), which is comparable to, or faster than, methods specialized for each task (see Table 1 in the Supplementary). Moreover, combining it with per-task heads allows us to train a relatively small number of parameters for new tasks, which we show by freezing the system and adding a head for motion-based segmentation and camera pose estimation (marked in green in Figures 1 and 2). Lastly, but perhaps most importantly, breaking the architecture into a general VideoMAE and per-task heads offers a mechanism to solve both dense and sparse tasks with a unified model (Figure 1). We emphasize that our main target is a general, multi-task framework, rather than the performance on any specific tasks. Nonetheless, our model performs competitively on the individual tasks, often on par with state-of-the-art methods. This is remarkable because the baselines we compare against are task-specific, carefully designed for, and specialized to excel at their respective tasks. Finally, VideoMAEs are already used as encoders for VLMs [76, 38], and we speculate that training them to reason about low-level 4D perception may impart those capabilities to the downstream VLMs they may be used with, though validating this is outside the scope of our paper. In summary:
  • We show that priors from video representation learning can be leveraged for low-level 4D perception tasks with strong generalization capabilities.
  • We present a feedforward video models that adopts a general-purpose architecture for multiple 4D perception tasks jointly—including both dense and sparse ones—and demonstrate its effectiveness on in-the-wild video inputs.
  • We combine the general-purpose architecture with a novel memory mechanism to support long-range tracking to tackle the limited context of the video encoder.
  • Our model demonstrates strong performance, solving multiple tasks jointly with results on par with state-of-the-art, within timeframes typical of single-task approaches.
@@ 2.1 Foundation models for vision perception
Self-supervised pre-training of large models on huge unlabeled data has shown great success. Among them, Vision Transformers (ViT) [33] pre-trained with masked autoencoding (MAE) [23] have become a common choice to fine-tune for many vision tasks, Segment Anything (SAM) [32] being a notable example. Several approaches have been proposed for spatio-temporal representation learning for videos: VideoMAE [64, 68], MAE-ST [16], V-JEPA [1], etc. We adopt VideoMAEv2 [68], which introduces a dual masking strategy that allows them to efficiently scale up the model to a billion parameters and to effectively leverage priors from large data. Their representation is effective for action recognition, and more recently is combined with language models [76, 38]. We show how to leverage them for low-level 4D perception tasks and achieve results competitive with specialized SOTA methods.
@@ 2.2 Dense Prediction Tasks
Depth estimation. Stereo depth estimation approaches [18, 19, 52, 87, 25, 50] and single image depth estimation approaches [15, 46, 47, 85, 86] have made significant progress. We focus on estimating depth from videos of dynamic scenes, for which several solutions exist [40, 34, 59, 75]. Pioneered by DUSt3R [71], an emerging paradign estimates pointmaps instead of depthmaps. Several recent approaches [89, 37, 70, 67] follow this paradigm for joint depth and camera pose estimation. To tackle the data scarcity problem, the state-of-the-art approaches leverage priors from foundation models: DepthAnything [85, 86] uses DINOv2 [43], DUSt3R [71] uses cross-view completion pre-training [77], and Marigold [31] and DepthCrafter [24] use image and video diffusion priors respectively. We benefit from large-scale pre-training by using a video encoder pre-trained with MAE.
@@
Camera pose estimation. Classical structure-from-motion solutions [51, 56] work robustly and accurately in constrained settings, while learning-based solutions [73, 61, 39, 71, 65, 89, 70, 67] address challenging scenarios like limited image observations, dynamic and texture-less regions, etc. We cast camera pose estimation as a dense task by representing cameras as dense bundles of rays [21, 90] using 6-D Plücker coordinates [45] and estimating them using our dense head.
@@ 2.3 Sparse Prediction Tasks
Tracking Any Point (TAP) in a video has many applications [79, 2, 36, 66, 9]. Particle Video [22], PIPs [4] and TAP-Net [11] lay the initial foundation by adopting ideas from optical flow approaches. On the other hand, OmniMotion [69] optimizes a volumetric representation for each video to solve this task and is time-consuming. TAPIR [12] introduces the idea of coarse-to-fine track estimation, BootsTAPIR [13] further improves it by adopting a self-supervised learning, and CoTracker [30] proposes to jointly track multiple points to leverage spatial correlations. SpaTracker [80] is a feedforward method for 3D point tracking that lifts pixels to 3D using input depth maps and applies CoTracker’s tracking formulation in 3D. Concurrent works, TAPIP3D [88] and SpatialTrackerV2 [81], adopt a similar approach but use posed RGBD inputs; notably, SpatialTrackerV2 jointly optimizes point tracks and initial camera poses via iterative bundle adjustment, which is inefficient for real-time settings. Seurat [10] shows that dense 2D tracks alone contain enough information to be lifted to 3D. In contrast to these prior works, which often rely on specialized architectures and additional inputs such as depth maps or 2D point tracks, we focus on efficiently unifying 2D and 3D point tracking with other tasks, paving the way for general purpose low-level 4D perception for real-time scenarios.
@@ 3.2 Dense Prediction Heads
For videos longer than $T$ frames, we run inference with stride $T/2$ and enforce consistency in the overlapping frames between temporal windows. For depth, we align the predictions with an affine transformation. This strategy has no effect on the individual windows for relative depth, but can greatly improve long-term temporal consistency. For optical flow and motion-based segmentation, we simply overwrite the predictions in the overlap between windows. For pose estimation, we align camera trajectories across overlapping windows using 3D similarity transformations between their uplifted point clouds.
@@ 3.3.2 A Memory Mechanism for Long Videos
To track beyond a single window of length $T$ frames, we propose an online strategy. A naïve approach would be to chain tracks across windows. Given two consecutive and overlapping windows, and a 2D track estimated in the first one, we can use a point on the track in the temporal overlap between the two windows as the query for tracking in the second one. To pick a good point to chain the tracks, we can select the one with highest visibility score. However, this solution is brittle for two reasons. First, a tracked point may not be visible in the overlap between the windows. To tackle this problem, inspired by Karaev et al[30], we introduce a track-feature token $\mathcal{F}$ that is passed to subsequent windows as an additional prompt (see Figure 3). However, unlike Karaev et al., we do not initialize it explicitly with the local appearance around the query point, so the two-way attention head is free to capture the most useful information to track through occlusions. Second, the naïve solution described above does not allow the system to reason across temporal windows, which makes it prone to drifting or to losing tracks. The track-feature tokens help, but to provide even more cross-window information, we pass the video tokens decoded by the two-way attention stage of the current window to the next, as shown in Figure 3. We achieve this by projecting the decoded video tokens in the overlapping region via a linear layer, and by adding them to the corresponding incoming video tokens to the two-way attention stage in the next window. Our memory strategy based on these two mechanisms is critical to allow proper reasoning across temporal windows as shown by the comparison in ablation study in Section 5.4.
@@ 4 Implementation
Training datasets. We use the video encoder from Wang et al[68], which is pre-trained on 1.35M video clips for masked auto-encoding. To fine-tune our model, we use four synthetic datasets covering various types of annotations, and rely on the priors from the pre-trained video encoder for generalization. We use Kubric [20], a synthetic dataset in which multiple objects interact, annotated with ground truth (GT) depth, flow, and 2D/3D point tracking. To include videos with long 3D trajectories, we add PointOdyssey [91] and DynamicReplica [29]. Both are synthetic datasets with animated characters in mostly indoor scenes. Both have depth GT and DynamicReplica also offers optical flow GT. To further increase scene diversity, we also include TartanAir [74], which provides GT for flow and depth.
@@ 5 Experiments
For each task included in our main model, we compare against a few recent SOTA methods on standard datasets. Our focus is a general, multi-task framework rather than achieving SOTA on any one task. Overall, we demonstrate competitive performance across all tasks, with computation times that are comparable to or faster than those of single-task approaches. We also show that L4P can be extended to new tasks without degrading existing performance.
@@
AbsRel $\downarrow$ $\delta_{1}$ $\uparrow$ AbsRel $\downarrow$ $\delta_{1}$ $\uparrow$ AbsRel $\downarrow$ $\delta_{1}$ $\uparrow$ AbsRel $\downarrow$ $\delta_{1}$ $\uparrow$ AbsRel $\downarrow$ $\delta_{1}$ $\uparrow$
@@
Table 1: Depth estimation results. We compare our full model with single-image (Row 1-3) and video depth estimation baselines (Row 4-7). On video datasets (all except NYUv2), our model consistently performs better than DepthCrafter, the closest competition, and by a large margin on ScanNet and KITTI. Frames indicate average video length for each dataset.
Figure 4: Comparisons with SOTA methods. We run inference on videos but show one frame here. The purple masks show GT motion segmentation.
@@ 5.1 Video Depth Estimation
We follow DepthCrafter [24] and evaluate video depth estimation on a collection of five datasets. We do not use any of the datasets for training our models or the baselines to better understand their generalization abilities. When estimating depth for videos longer than $T=16$ frames, we follow the online inference and alignment strategies explained in Section 3.2. There is an inherent scale-ambiguity in the estimated depthmaps. We follow the common practice of aligning linearly the estimation with the GT before calculating evaluation metrics. The alignment is done for all the frames at once, and is carried out in disparity space via least-square fitting. For comparison on single image datasets, we repeat the single frame 16 times to compute our estimations. We report two metrics: AbsRel ($\text{mean}(|\hat{\mathbf{d}}-\mathbf{d}|/\mathbf{d}))$ and $\delta_{1}$ (ratio of pixels satisfying $\max(\mathbf{d}/\hat{\mathbf{d}},\hat{\mathbf{d}}/\mathbf{d})<1.25$), where $\mathbf{d}$ represents GT, and $\hat{\mathbf{d}}$ is depth estimation after alignment. We upsample our estimations from $224\times 224$ to each dataset’s original resolution for evaluation. We consider video approaches including NVDS [75], ChronoDepth [54], DepthCrafter [24], MonST3R [89], as well as single-image ones, including Marigold [31] and DepthAnything [85, 86].
Our results show strong performance against both single-image and video depth approaches on the four video datasets (Table 1). Since L4P is a video approach, applying it on single images from NYUv2 does not provide the necessary temporal context for it to perform well. DepthCrafter and MonST3R similarly suffer on NYUv2. Figure 4 (first row) shows a qualitative comparison. L4P produces a level of details comparable to that of diffusion models such as DepthCrafter, while generally capturing more accurate relative scales. It is also worth noting that our approach performs competitively on KITTI, despite the fine-tuning not including driving scenarios.
Spring Sintel Virtual KITTI
$EPE\downarrow$ $1\uparrow EPEEPE\downarrow$ $1\uparrow EPEEPE\downarrow$ $EPE<1\uparrow$
RAFT* [60] $0.13$ $98.4$ $1.31$ $85.1$ $1.13$ $77.5$
MemFlow [14] $0.11$ $\mathbf{98.7}$ $\mathbf{1.03}$ $\mathbf{87.2}$ $0.72$ $85.4$
L4P $\mathbf{0.09}$ $\mathbf{98.7}$ $1.06$ $84.6$ $\mathbf{0.63}$ $\mathbf{86.6}$
Table 2: Optical flow estimation results. We compare with SOTA flow methods: RAFT (two-frame, marked with *) and MemFlow (multi-frame). We evaluate cross-dataset generalization at 224x224 resolution. Our approach is competitive at this resolution, despite using generic architecture.